Bob is two times older than Tim. 25 years ago Bob was three times older than tim. How old are they?

Answer by Alan3354(69443)

Bob is two times older than Tim. 25 years ago Bob was three times older than tim. How old are they?
----------
B = 3T
B - 25 = 4*(T - 25)
------
3T - 25 = 4T - 100
T = 75
Tim is 75
Bob is 225
==================
They must be tortoises.